OpenStudy (anonymous):

Solve: 4(2x - 1) - 3 = 9(x - 4) -29 32 13 29

OpenStudy (solomonzelman):

\[4(2x - 1) - 3 = 9(x - 4)\]\[8x - 4 - 3 = 9x - 36\]\[8x - 7 = 9x - 36\]\[ -7 + 36 = 9x -8x\]\[x=29\]

OpenStudy (anonymous):

Thanks!

OpenStudy (solomonzelman):

Anytime!